Factor the polynomial
x^3+9x^2+26x+24

Answer by lwsshak3(6505)

Factor the polynomial
x^3+9x^2+26x+24
Use the rational roots theorem to solve:
...0..|.....1......9.......26......24
...1..|.....1......10.....36......60 (1 is upper limit, all numbers>0)
=================
.-1..|.....1......8.......18.......6
.-2..|.....1......7.......12.......0 (-2 is a root)
(x+2)(x^2+7x+12)
=(x+2)(x+4)(x+3)
x^3+9x^2+26x+24=(x+2)(x+4)(x+3)